OpenStudy (anonymous):

1) A two-pen corral is to be built. The outline of the corral forms two identical adjoining rectangles. If there is a 120 ft. of fencing available, what dimensions of the corral will maximize the enclosed area?

OpenStudy (anonymous):

@electrokid can u help? @Mertsj

OpenStudy (anonymous):

if the dimensions of one pen are x by y, then the area is xy, y is the dimension in common. area of both A = 2xy perimeter = 120 = 4x+3y combine the two equations to get A as a function of x or y 120 = 4x+3y 4x = 120–3y x = 30 –(3/4)y A = y(30 –(3/4)y)= 30y –(3/4)y² differentiate and set equal to 0 to get max/min A' = 30 – (3/4)2y = 0 (3/2)y = 30 y = 20 x = 30 –(3/4)y = 30 –(3/4)20 = 15 pens are 20 x 15, and overall is 30 x 20

OpenStudy (anonymous):

huh?

OpenStudy (mertsj):

The solution provided is exactly right. What class are you taking that you have no understanding of the solution?

OpenStudy (anonymous):

Noi thought he would explain it but okay